OpenStudy (anonymous):

Create your own binomial expression with a radical in the second term. Part 1: Identify its conjugate and explain, in complete sentences, why it is the conjugate. (1 point) Part 2: Multiply your original binomial expression and its conjugate. What happened to the radicals and why? (1 point)

OpenStudy (anonymous):

still nobody gave you answer?

OpenStudy (anonymous):

\[(a+\sqrt{b})\] conjugate: \[(a-\sqrt{b})\] multiply: \[(a-\sqrt{b})(a+\sqrt{b})= a ^{2} -a \sqrt{b} + a \sqrt{b} -\sqrt{b}^{2}= a ^{2} -b\]